Subject: 2.5m obslog 20080330 (MJD 54556) Date: Mon, 31 Mar 2008 05:57:40 -0600 (MDT) From: hjbrew@apo.nmsu.edu (Howard Brewington) APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday March 30, 2008 (MJD 54556) ---=== OBSERVING TEAM ===--- Shannon Watters (observer) Howard Brewington (observer) Dmitry Bizyaev (dayshift) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Declared three plates DONE, viz., 2917, 2791, and 2793. Added signal to two plates, viz., 2960 and 2874. We completed U-1 linearity MSP; run # 7404. The kick-plate on the laptop-side of the doghouse drags on the floor when the scope points to the southeast. The alt-locking pin was sticking after most cart changes. The dial tone of the phone at the laptop (line #120) was working intermittently. The power to the manipulator arm had to be cycled several times during the night because the arm would stop moving while placing a cartridge in the bay or on the Lindy cart. The dust averaged about 2.4K for the night, so the mirrors need to be checked. The spectro dewar needs to be changed. ---=== OBSERVING LOG ===--- Afternoon: ---------- ntp_check ok iccSoundOff ok imager bias ok spectro bias ok idlspec2d> apo_plotbias, 51452 Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00051452.fit -6.3 -4.2 -3.1 2.1 7.3 9.0 10.5 sdR-b2-00051452.fit -7.9 -5.8 -4.6 0.9 7.2 8.9 11.0 sdR-r1-00051452.fit -6.7 -5.6 -3.6 1.4 6.4 8.4 10.4 sdR-r2-00051452.fit -7.7 -6.4 -4.6 0.9 6.2 8.0 10.1 no data safe to remove from APO Collimation is good for ~+8C. Ventilation started at 00:10Z. Cartridge 5 is mounted Night: ------ 01:15Z Enclosure off, topped-off spectro secondary dewars, checked top/bottom balance, did a test slew, checked fiducial crossings for each axis, started the DIMM. 01:20Z Sunset 01:50Z Cart 05, plate 2960 (boaTest187). The field was found with an fk5 and spiralPattern. Seeing was ~1.5" as reported by the guider. The DIMM was not providing any useful data because of high winds. One exposure was completed; see below. 02:26 51458 10.0 flat 02:29 51459 2.0 arc 02:44Z Twilight ended 02:59 51460 1501.0 target 03:12Z Cart 03, plate 2874 (segFaint170); TCC timed out as we slewed to the field. An fk5 was used to find the field. Two exposure were completed; see below. 03:26 51461 10.0 flat 03:29 51462 2.0 arc 03:58 51463 1500.0 target 04:27 51464 1502.0 target 04:30 51465 10.0 flat 04:32 51466 2.0 arc 04:40Z Cart 06, plate 2917 (segFaint180); the TCC timed out again during the slew to the field. Guide stars on goToField. However, we were near the wind-closure limit, and we have clouds forming over the site. 05:00Z Enclosure on, doors/louvers closed, and fans off because of dust and high winds, i.e., gusts = 42 MPH. 05:45Z Enclosure off, fans on, louvers open as wind and clouds decreased. We did not reopen the DIMM because of wind gusts near 30 MPH. 05:51Z Cart 06, plate 2917 (segFaint180); Guide stars on goToField. Seeing was about 1.8" as reported by the guider. Three exposures were completed, and the plate declared DONE. See below. 05:59 51467 10.0 flat 06:02 51468 2.0 arc 06:25 51469 1200.0 target 06:48 51470 1201.0 target 07:06 51471 902.0 target 07:10 51472 10.0 flat 07:13 51473 2.0 arc MJD b1 r1 b2 r2 ------------------------------ 54555 78.1 36.0 86.9 46.5 54556 33.2 18.1 37.9 22.6 ------------------------------ Total 111.3 54.1 124.8 69.1 07:24Z Cart 02, plate 2791 (chunk94). An fk5 was used to find the field. Seeing was about 1.5" as reported by the guider. Three exposures were completed; the plate was declared DONE. 07:28 51474 10.0 flat 07:30 51475 2.0 arc 07:50 51476 900.0 target 08:02 51477 601.0 target 08:17 51478 702.0 target 08:23Z Cart 09, plate 2792 (chunk165). Guide stars found on goToField. The seeing was about 2.0" as reported by the guider. This plate was declared DONE after three exposures; see below. 08:37 51479 10.0 flat 08:40 51480 2.0 arc 09:01 51481 900.0 target 09:29 51482 1501.0 target 09:52 51483 1202.0 target 09:55Z Mounted cart 05 10:00Z Enclosure on, doors/louvers closed, fans off because of dust ( 3000+ ) and humidity ( >50% ). Integrated dust level reached ~20K dust hours for the night. 10:10Z We setup to do U-1 linearity MSP. The run number was 7404. 10:50Z Dust = 4371; humidity = 52.1% 11:33Z Morning twilight 11:54Z endNight finished and was happy ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7403 22:57Z 23:13Z 100 O -117.75 -113.87 39 ignore 7405 11:42Z 11:45Z 100 O 73.92 74.61 19 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 51451, plate -9999 ------- 22:33 51451 0.0 bias 23:01 51452 0.0 bias ----- sequence 51453, plate 2177 ------- 23:21 51453 10.0 flat 23:24 51454 2.0 arc 23:32 51455 3.0 arc 23:34 51456 3.0 arc ----- sequence 51457, plate 0 ------- 01:09 51457 0.0 bias ----- sequence 51458, plate 2960 ------- 02:26 51458 10.0 flat 02:29 51459 2.0 arc 02:59 51460 1501.0 target ----- sequence 51461, plate 2874 ------- 03:26 51461 10.0 flat 03:29 51462 2.0 arc 03:58 51463 1500.0 target 04:27 51464 1502.0 target 04:30 51465 10.0 flat 04:32 51466 2.0 arc ----- sequence 51467, plate 2917 ------- 05:59 51467 10.0 flat 06:02 51468 2.0 arc 06:25 51469 1200.0 target 06:48 51470 1201.0 target 07:06 51471 902.0 target 07:10 51472 10.0 flat 07:13 51473 2.0 arc ----- sequence 51474, plate 2791 ------- 07:28 51474 10.0 flat 07:30 51475 2.0 arc 07:50 51476 900.0 target 08:02 51477 601.0 target 08:17 51478 702.0 target ----- sequence 51479, plate 2792 ------- 08:37 51479 10.0 flat 08:40 51480 2.0 arc 09:01 51481 900.0 target 09:29 51482 1501.0 target 09:52 51483 1202.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:35Z 5 2960 88.05 0.0001 82.59 0.0074 263.35 0.0336 1.000040 03:31Z 3 2874 41.52 0.0023 34.93 0.0045 214.73 0.0455 1.000080 06:06Z 6 2917 11.59 0.0008 56.66 0.0046 189.69 0.0090 1.000140 07:35Z 2 2791 68.91 0.0011 62.21 0.0061 236.39 0.0223 1.000140 09:23Z 9 2792 10.37 -0.0019 76.21 0.0036 189.18 0.0426 1.000120 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 00:51Z 5 0 1.00000 0 1257 0 121 30.0 10.9 28 246 02:35Z 5 2960 1.00004 -395 889 100 88 82.6 9.3 24 265 gdr 1.5 03:32Z 3 2874 1.00008 -790 682 100 41 35.0 8.7 26 253 gdr 2.0 06:06Z 6 2917 1.00014 -1383 117 100 12 56.6 6.3 21 244 gdr 1.8 07:35Z 2 2791 1.00014 -1383 203 200 69 62.2 5.5 19 267 gdr 1.5 09:22Z 9 2792 1.00015 -1481 71 175 11 76.2 4.2 25 268 gdr 2.0 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:36Z 57 15 24 205 (SSW) 1807 - 23:07Z 54 11 22 261 (W) 1605 - 23:37Z 54 11 19 234 (SW) 2441 - 00:07Z 55 12 27 239 (WSW) 1882 - 00:39Z 52 10 24 263 (W) 1943 - 01:09Z 50 10 25 241 (WSW) 1481 - 01:41Z 48 10 26 255 (WSW) 2492 - 02:12Z 49 10 30 245 (WSW) 2167 - 02:43Z 48 11 21 264 (W) 2178 - 03:14Z 47 13 27 250 (WSW) 2026 - 03:45Z 47 14 25 255 (WSW) 2151 - 04:15Z 47 15 29 261 (W) 2111 - 04:48Z 46 17 28 251 (WSW) 2301 - 05:18Z 45 16 24 253 (WSW) 2382 - 05:51Z 43 15 24 249 (WSW) 2350 - 06:21Z 42 15 18 251 (WSW) 2444 - 06:53Z 42 15 14 265 (W) 2389 - 07:23Z 41 16 23 270 (W) 2334 - 07:55Z 41 18 14 268 (W) 2342 - 08:25Z 40 20 18 278 (W) 2524 - 08:56Z 40 21 13 278 (W) 2578 - 09:27Z 39 22 22 276 (W) 3153 - 09:58Z 38 22 22 277 (W) 3176 - 10:29Z 38 22 10 265 (W) 2804 - 11:00Z 37 22 20 265 (W) 4351 - 11:32Z 36 23 24 260 (W) 2883 - ---=== TELESCOPE STATUS ===--- 00:10Z Doors and louvers open, fans on. 01:15Z Enclosure off. 05:00Z Enclosure on, doors and louvers closed, fans off. 05:45Z Enclosure off, fans on, louvers open 10:00Z Enclosure on, louvers/doors closed, fans off No interlocks bypassed. Status at: 12:00Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 05 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 77lbs, 19 psi Imager 170lbs, 20 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_39 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0170 0.0200 Axial B -0.0070 0.0000 Axial C -0.0270 -0.0230 Trans D -0.0010 0.0120 Lateral E 10.0838 10.0584 Lateral F 11.6459 11.6205 GALILS Commanded: 4600. -2000. 2650. -12450. -6500. -6550. Actual: 4600. -2000. 2650. -12439. -6477. -6546. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 100.00 -200.00 PriOrient: 0.00 0.00 0.00 99.59 -200.69 SECONDARY: ---------- Focus: 0.00 Air Temp.: 12.0 Alt.: 30.000004 MIGS TONIGHT NOMINAL Axial A -0.0350 -0.0440 Axial B -0.0490 -0.0570 Axial C -0.0600 -0.0670 Trans D -0.0170 -0.0110 GALILS Commanded: 1627472. 1653110. 1470900. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.98 59.98 -4.01 -19.92 119.30 ---=== PROBLEMS IN DETAIL ===--- Collimation (with cartridge 4) sp1> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 8.22429 tpm_TM_TelairAvg.VALA 9.88 sp1> mechSend s spMechVersion v2_9_0 Bootup 870616738 SpectroID 1 SlitID 36 Air On Shutter_open_sensor Off Shutter_closed_sensor On Left_open_sensor Off Left_closed_sensor On Right_open_sensor Off Right_closed_sensor On Coll_motor_A -52 01P-0000000052 Coll_motor_B 117 02P0000000117 Coll_motor_C -169 01P-0000000169 Coll_motor_A_status 132 01S0132 Coll_motor_B_status 133 02S0133 Coll_motor_C_status 133 01S0133 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 3 Last_open_transit.time 700 Last_close_transit.time 545 Humid_Hartmann 14.2 0xFDFF Humid_Cen_Optics 12.3 0xFE18 Temp_Median 10.7 Temp_Hartmann_Top 10.3 (0xFE4F) Temp_Red_Cam_Bot 10.7 (0xFE3F) Temp_Red_Cam_Top 10.9 (0xFE37) Temp_Blue_Cam_Bot 11.2 (0xFE2B) Temp_Blue_Cam_Top 10.7 (0xFE3F) AD 28160 1294 sp2> mechSend s spMechVersion v2_9_0 Bootup 870616738 SpectroID 2 SlitID 4 Air On Shutter_open_sensor Off Shutter_closed_sensor On Left_open_sensor Off Left_closed_sensor On Right_open_sensor Off Right_closed_sensor On Coll_motor_A 2619 01P0000002619 Coll_motor_B 2132 02P0000002132 Coll_motor_C 2053 01P0000002053 Coll_motor_A_status 133 01S0133 Coll_motor_B_status 132 02S0132 Coll_motor_C_status 133 01S0133 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 3 Last_open_transit.time 710 Last_close_transit.time 625 Humid_Hartmann -24.9 0x0 Humid_Cen_Optics 12.6 0xFE14 Temp_Median 8.6 Temp_Hartmann_Top 9.3 (0xFE7C) Temp_Red_Cam_Bot 9.4 (0xFE78) Temp_Red_Cam_Top 8.4 (0xFEA0) Temp_Blue_Cam_Bot 8.9 (0xFE8C) Temp_Blue_Cam_Top 8.7 (0xFE94) AD 22335 1094 IDL 5.5a> collimate, 51455 Reading /data/spectro/54556/sdR-b1-00051455.fit Reading /data/spectro/54556/sdR-b1-00051456.fit COLLIMATE: Min offset = -0.87 pix COLLIMATE: Max offset = 0.64 pix COLLIMATE: RMS across CCD = 0.32 pix COLLIMATE: Mean offset = -0.06 pix COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 2.3 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54556-b1-00051455.log PostScript file = Collimate-54556-b1-00051455.ps Reading /data/spectro/54556/sdR-b2-00051455.fit Reading /data/spectro/54556/sdR-b2-00051456.fit COLLIMATE: Min offset = -0.34 pix COLLIMATE: Max offset = 0.16 pix COLLIMATE: RMS across CCD = 0.13 pix COLLIMATE: Mean offset = -0.06 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 2.1 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54556-b2-00051455.log PostScript file = Collimate-54556-b2-00051455.ps Reading /data/spectro/54556/sdR-r1-00051455.fit Reading /data/spectro/54556/sdR-r1-00051456.fit COLLIMATE: Min offset = -0.56 pix COLLIMATE: Max offset = 0.65 pix COLLIMATE: RMS across CCD = 0.32 pix COLLIMATE: Mean offset = -0.03 pix COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 308 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 308" Log file = Collimate-54556-r1-00051455.log PostScript file = Collimate-54556-r1-00051455.ps Reading /data/spectro/54556/sdR-r2-00051455.fit Reading /data/spectro/54556/sdR-r2-00051456.fit COLLIMATE: Min offset = -0.81 pix COLLIMATE: Max offset = 0.20 pix COLLIMATE: RMS across CCD = 0.26 pix COLLIMATE: Mean offset = -0.08 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 693 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 693" Log file = Collimate-54556-r2-00051455.log PostScript file = Collimate-54556-r2-00051455.ps Collimation should be good for ~+8C ambient -------------